Solve the following pDEs
A). d²u÷dx²= 8xy²+1
B). d²u÷dxy- du÷dx=6xe^x
1
Expert's answer
2019-06-18T10:48:48-0400

A)

"\frac{\partial^2u}{\partial x^2}= 8xy^2+1."


Integrate each side of the equation by x. Antiderivatives will differ by an arbitrary function C1(y):


"\frac{\partial u}{\partial x}= 4x^2y^2+x+C_1(y)."

Repeat this:


"u = \frac{4}{3}x^3y^2+\frac{x^2}{2}+C_1(y)x +C_2(y)."

C2(y) is an arbitrary function.

B)

"\frac{\partial ^2u}{\partial x \partial y} -\frac{\partial u}{\partial x}=6xe^x."


Integrate each side of the equation by x. Antiderivatives will differ by an arbitrary function C1(y):


"\frac{\partial u}{\partial y} - u = 6(x-1)e^x+C_1(y)."

This is linear ordinary differential linear equation by y. Then by the well-known formula

[https://en.wikipedia.org/wiki/Linear_differential_equation#First-order_equation_with_variable_coefficients]


"u = e^y\left(\int\limits_{y_0}^{y}{\Bigl(6(x-1)e^x + C_1(t)\Bigr)e^{-t}dt} + C_2(x) \right),"

where y0 is any preselected number (let y0=0 for example), C2(x) is an arbitrary function of x.

Then


"u = \left(6(x-1)e^x\int\limits_{y_0}^{y}{e^{-t}dt} + \int\limits_{y_0}^{y}{C_1(t)e^{-t}dt} + C_2(x)\right) e^y."


Now let the integrals not depend on x: this could be taken into account in arbitrary C2(x). Then consider a new arbitrary function C3(y) based on arbitrary function C1(y):

"C_3(y) = 
e^y\int\limits_{y_0}^{y}{C_1(t)e^{-t}dt},"


and a new arbitrary function C4(x) based on arbitrary function C2(x):

"\int\limits_{0}^{y}{e^{-t}dt} = -e^{-y} + 1,""C_4(x) = 6(x-1)e^x + C_2(x)."


Then the expression is simplified:


"u = C_4(x)e^y + C_3(y) - 6(x-1)e^x."